An aqueous magnesium chloride solution having a density of 1.049 g/cm3 is found to have an osmotic pressure of 17.0 atm at 25°C. Find the temperature at which this solution boils.
Answer #1

C- Os moléc Breleme cricRT - O C= Concertoweon Comel LT=Teruperature iu Kewin R= pas Constanta 0.00206 atab not é = Vant H

we know foon boiling - boênut elevation I oth = ex Kb x m - ② atb= To (solution) - Tb (bure solvent) Tocsolution) = oth + Tb